# Integral representations for the joint survival functions of the cumulated components of multinomial random vectors

### v1

@@ 3 Proofs
$=\exp\left\{\Delta_{N}+N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\},$
where recall $\Delta_{N}=\ln\{(N+d)!/(N!N^{d})\}+\Lambda_{N}-(1/2)\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon})$ from (4). Putting this last equation in (2) yields
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left[N\gamma^{\star}(\boldsymbol{s})\right]N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
@@
For all $\boldsymbol{s}\in\mathcal{S}_{d}$, define
@@
For all $\boldsymbol{k}\in\mathbb{N}_{0}^{d}\cap n\mathcal{S}_{d}$, the integral representation of the joint survival function of the cumulated components of any multinomial random vector derived in (2) can be rewritten as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])=\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\int_{\mathcal{R}_{d}}\exp\left\{NH(\boldsymbol{s})\right\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$ (13)
@@
where $(12m+1)^{-1}\leq\lambda_{m}\leq(12m)^{-1}$. Since $\sum_{i=1}^{d+1}J_{i}=N$ and $\Lambda_{N}=\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{i}$, one deduces that
$\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}=\frac{\exp(\Lambda_{N})}{\sqrt{(2\pi N)^{d}\prod_{i=1}^{d+1}\{p_{i}(1+\varepsilon_{i})\}}}\exp\left\{-NH(\boldsymbol{J}/N)\right\},$
@@
$\qquad=\frac{(N+d)!}{N!N^{d}}\times\frac{N^{d/2}\exp(\Lambda_{N})}{\sqrt{(2\pi)^{d}\prod_{i=1}^{d+1}\{p_{i}(1+\varepsilon_{i})\}}}\int_{\mathcal{R}_{d}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{J}/N)\right\}\right]\,\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
@@
[…]
@@
where recall $\widetilde{\boldsymbol{\varepsilon}}=(p_{1}\varepsilon_{1},\ldots,p_{d}\varepsilon_{d})^{\top}$, $\widetilde{\varepsilon}_{d+1}=p_{d+1}\varepsilon_{d+1}$, and $\Sigma_{\boldsymbol{p}}^{-1}=\{p_{i}^{-1}\mathds{1}(i=j)+p_{d+1}^{-1}\}_{1\leq i,j\leq d}$. Recall from (4) that
@@
$\qquad=\frac{N^{d/2}\exp(\Delta_{N})}{\sqrt{(2\pi)^{d}\det(\Sigma_{\boldsymbol{p}})}}\int_{\mathcal{R}_{d}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{p})\right\}-\frac{N}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}\right]\,\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
@@
$=\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})$
$\qquad-\frac{1}{2}\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}}$
@@
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{J}/N-\boldsymbol{s})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}$
$\qquad=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
@@
The choice of decomposition in (15) can be explained as follows. The Hessian matrix of $H$ at $\boldsymbol{s}$ is equal to
$\left(-\frac{J_{i}}{Ns_{i}^{2}}\mathds{1}(i=j)-\frac{J_{d+1}}{Ns_{d+1}^{2}}\right)_{1\leq i,j\leq d},$

### v2

@@ 4 Proofs
[…]
where we recall that $\Delta_{N}=\ln\{(N+d)!/(N!N^{d})\}+\Lambda_{N}-(1/2)\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon})$ from (6). Putting this last equation in (4) yields
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~\forall i\in[d])=e^{\Delta_{N}}\int_{\mathcal{R}_{d}^{\circ}}\exp\left[N\gamma^{\star}(\boldsymbol{s})\right]N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$
@@
For all $\boldsymbol{s}\in\mathrm{Int}(\mathcal{S}_{d})$, with $s_{d+1}=1-\|\boldsymbol{s}\|_{1}$, define
@@
For all $\boldsymbol{k}\in\mathcal{K}_{n,d}$, the integral representation of the joint survival function of the cumulative sums of the components of any multinomial random vector derived in (4) can be rewritten as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~\forall i\in[d])=\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\int_{\mathcal{R}_{d}^{\circ}}\exp\left\{NH(\boldsymbol{s})\right\}\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$ (15)
@@
where $(12m+1)^{-1}\leq\lambda_{m}\leq(12m)^{-1}$. Since $\sum_{i=1}^{d+1}J_{i}=N$ and $\Lambda_{N}=\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{J_{i}}$, one deduces that
$\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}=\frac{\exp(\Lambda_{N})}{\sqrt{(2\pi N)^{d}\prod_{i=1}^{d+1}\{p_{i}(1+\varepsilon_{i})\}}}\exp\left\{-NH(\boldsymbol{J}/N)\right\}.$
@@
$\qquad=\frac{(N+d)!}{N!N^{d}}\times\frac{N^{d/2}\exp(\Lambda_{N})}{\sqrt{(2\pi)^{d}\prod_{i=1}^{d+1}\{p_{i}(1+\varepsilon_{i})\}}}\int_{\mathcal{R}_{d}^{\circ}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{J}/N)\right\}\right]\,\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$
@@
$\int_{\mathcal{R}_{d}^{\circ}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{p})\right\}+N\left\{H(\boldsymbol{p})-H(\boldsymbol{J}/N)\right\}\right]\,\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$ (17)
@@
where we recall that $\widetilde{\boldsymbol{\varepsilon}}=(p_{1}\varepsilon_{1},\ldots,p_{d}\varepsilon_{d})^{\top}$, $\widetilde{\varepsilon}_{d+1}=p_{d+1}\varepsilon_{d+1}$, and $\Sigma_{\boldsymbol{p}}^{-1}=\{p_{i}^{-1}\mathds{1}\{i=j\}+p_{d+1}^{-1}\}_{1\leq i,j\leq d}$. Recall from (6) that
@@
$\qquad=\frac{N^{d/2}\exp(\Delta_{N})}{\sqrt{(2\pi)^{d}\det(\Sigma_{\boldsymbol{p}})}}\int_{\mathcal{R}_{d}^{\circ}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{p})\right\}-\frac{N}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}\right]\,\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$
@@
$=\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}}$
@@
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~\forall i\in[d])$ $=e^{\Delta_{N}}\int_{\mathcal{R}_{d}^{\circ}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{J}/N-\boldsymbol{s})\}\thinspace\mathrm{d}\thinspace\boldsymbol{s}$
$=e^{\Delta_{N}}\int_{\mathcal{R}_{d}^{\circ}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$
@@
The choice of decomposition in (17) can be explained as follows. The Hessian matrix of $H$ at $\boldsymbol{s}\in\mathrm{Int}(\mathcal{S}_{d})$ is equal to
$\left(-\frac{J_{i}}{Ns_{i}^{2}}\mathds{1}\{i=j\}-\frac{J_{d+1}}{Ns_{d+1}^{2}}\right)_{1\leq i,j\leq d},$